OpenStudy (anonymous):

lim x-> infinity ((\sqrt(x^2-16))-(\sqrt(x^2-1)))

OpenStudy (anonymous):

\[\lim x \rightarrow \infty (\sqrt{x^2+16}-\sqrt{x^2-1})\]

OpenStudy (anonymous):

@JimSacc HINT:- multiply and divide by x

OpenStudy (anonymous):

let me try that i think i know what you mean

OpenStudy (anonymous):

nah i don't know whats the next step

OpenStudy (anonymous):

\[x*\frac{ \sqrt{x^2+16}-\sqrt{x^2-1)} }{ x }\] x*\[x*(\sqrt{\frac{ x^2+16 }{ x^2 }} - \sqrt{\frac{ x^2-1 }{x^2 }})\]

OpenStudy (anonymous):

\[x*(\sqrt{1+\frac{ 16 }{ x^2 }} - \sqrt{1-\frac{ 1 }{ x^2 }})\] now apply the limits

OpenStudy (anonymous):

yeah so when it says x goes to infinity what number do i plug for x, i apologize for my complete ignorance

OpenStudy (anonymous):

u need to plug in infinity

OpenStudy (anonymous):

oh alright 0... haven't been studying appropriately at all this semester

OpenStudy (anonymous):

yeah 0 will be an answer

OpenStudy (anonymous):

@niksva, plugging \(\infty\) in directly will yield \(0\cdot\infty\), not \(0\). @JimSacc, multiply by conjugate/conjugate: \[\lim_{x\to\infty}\left(\sqrt{x^2-16}-\sqrt{x^2-1}\right)\cdot\frac{\sqrt{x^2-16}+\sqrt{x^2-1}}{\sqrt{x^2-16}+\sqrt{x^2-1}}\\ \lim_{x\to\infty}\frac{\left(x^2-16\right)-\left(x^2-1\right)}{\sqrt{x^2-16}+\sqrt{x^2-1}}\\ -15\lim_{x\to\infty}\frac{1}{\sqrt{x^2-16}+\sqrt{x^2-1}}\] Now the function properly approaches 0 as \(x\to\infty\).

OpenStudy (mathmale):

Which do you mean: 1) lim x-> infinity ((\sqrt(x^2-16))-(\sqrt(x^2-1))) or 2)\[\lim x \rightarrow \infty (\sqrt{x^2+16}-\sqrt{x^2-1})\] I ask this because in one expression you have Sqrt(x^2+16) and in the other you have Sqrt(x^2-16).

OpenStudy (anonymous):

@SithsAndGiggles any number multiplies with zero is always a zero

OpenStudy (anonymous):

That is not the case, in general. Consider \[\lim_{x\to\infty}e^xe^{-x}\] Plugging in directly yields \(0\cdot\infty\). However, \(e^{x}e^{-x}=1\), not \(0\).

OpenStudy (anonymous):

If the limit is as @mathmale indicated: \(\color{blue}{\text{Originally Posted by}}\) @SithsAndGiggles @JimSacc, multiply by conjugate/conjugate: \[\lim_{x\to\infty}\left(\sqrt{x^2\color{red}{+}16}-\sqrt{x^2-1}\right)\cdot\frac{\sqrt{x^2\color{red}{+}16}+\sqrt{x^2-1}}{\sqrt{x^2\color{red}{+}16}+\sqrt{x^2-1}}\\ \lim_{x\to\infty}\frac{\left(x^2\color{red}{+}16\right)-\left(x^2-1\right)}{\sqrt{x^2\color{red}{+}16}+\sqrt{x^2-1}}\\ \color{red}{17}\lim_{x\to\infty}\frac{1}{\sqrt{x^2\color{red}{+}16}+\sqrt{x^2-1}}\] Now the function properly approaches 0 as \(x\to\infty\). \(\color{blue}{\text{End of Quote}}\) The result is not affected.

OpenStudy (anonymous):

@SithsAndGiggles answer is always going to be zero any number multiplied with zero is always be zero , there might be some exceptions But i m 99 percent sure it is always going to be zero the case provided by you exp(x)*exp(-x) is always going to be equal to 1 according to my knowledge \[\exp(-x) = \frac{ 1 }{ \exp(x) }\]

OpenStudy (anonymous):

or one can use basic property \[\exp(x) * \exp(y) = \exp(x+y)\] after using this property i get 1 directly